# Confluent functions, Laguerre polynomials and their (generalized) bilinear integrals

### v1

@@ A.1 Definition
Definition A.1.
We say that a function $f$ on ]0,[]0$\infty$[] 0 , ∞ [ is integrable in the generalized sense if it is integrable on ]1,[]1$\infty$[] 1 , ∞ [ and there exists a finite set $\Omega\subset{\mathbb{C}}$ and complex coefficients $(f_{k})_{k\in\Omega}$ such that
$f-\sum_{k\in\Omega}f_{k}r^{k}$ (A.1)
is integrable on ]0,1[]0,1[] 0 , 1 [. The set of such functions is denoted ${\cal F}$. For $f\in{\cal F}$ we define
$\mathrm{gen}\int_{0}^{\infty}f(r){\mathrm{d}}r$ $\coloneqq\sum_{k\in\Omega\backslash\{-1\}}\frac{f_{k}}{k+1}+\int_{0}^{1}\Big{(}f(r)-\sum_{k\in\Omega}f_{k}r^{k}\Big{)}{\mathrm{d}}r+\int_{1}^{\infty}f(r){\mathrm{d}}r.$ (A.2)
Note that the set $\{k\in\Omega\,|\,\mathrm{Re}(k)\leq-1\}$ and the corresponding $f_{k}$ are uniquely determined by $f$. It is convenient to allow $k$ with $\mathrm{Re}(k)>-1$. The generalized integral of $f$ does not depend on the choice of $\Omega$.
@@
$\mathrm{gen}\int_{0}^{\infty}f(r){\mathrm{d}}r=\int_{0}^{\infty}f(r){\mathrm{d}}r\quad\text{for }f\in L^{1}]0,\infty[.$ (A.3)
@@
Definition A.2.
The generalized integral (A.2) is called anomalous if there exists $n\in{\mathbb{N}}$ such that $f_{-n}\neq 0$.
A.2 Integration by parts
Definition A.3.
We say that $F\in{\cal G}$ if $F$ is a measurable function on [0,[[0$\infty$[[ 0 , ∞ [, bounded on [1,[[1$\infty$[[ 1 , ∞ [ and there exists a finite set $\Theta\subset{\mathbb{C}}$ and complex coefficients $(F_{k})_{k\in\Theta}$ and $c\in{\mathbb{C}}$, such that
$F(r)-\sum_{k\in\Theta}F_{k}r^{k}-c\ln r$ (A.4)
has a limit as $r\to 0$. For $F\in{\cal G}$ we define the regular value of $F$ at 00:
$\mathop{\mathrm{rv}_{0}}F\coloneqq\lim_{r\searrow 0}\Big{(}F(r)-\sum_{k\in\Theta\backslash\{0\}}F_{k}r^{k}-c\ln r\Big{)}.$ (A.5)
Note that given $F$ the coefficients $F_{k}$ for $k\in\Theta$ such that $\mathrm{Re}k\leq 0,$ $k\neq 0$, as well as $c$, are uniquely defined. Hence (A.5) depends only on $F$.
Proposition A.4.
@@ A.2 Integration by parts
$\mathrm{gen}\int_{0}^{\infty}f(y){\mathrm{d}}y=-\mathop{\mathrm{rv}_{0}}F.$ (A.7)
@@
It is clear that $F$ is bounded on [1,[[1$\infty$[[ 1 , ∞ [ because $f\in{\cal F}$. Let $10<r<$ 1. Then
$\int_{r}^{\infty}f(y){\rm d}y$ $=\int_{r}^{1}\Big{(}f(y)-\sum_{k\in\Omega}f_{k}y^{k}\Big{)}{\mathrm{d}}y+\int_{1}^{\infty}f(y){\mathrm{d}}y$ (A.8)
$+\sum_{k\in\Omega\backslash\{-1\}}\frac{f_{k}}{k+1}(1-r^{k+1})-f_{-1}\ln r$
$=$ $\mathrm{gen}\int_{0}^{\infty}f(y){\mathrm{d}}y-\int_{0}^{r}\Big{(}f(y)-\sum_{k\in\Omega}f_{k}y^{k}\Big{)}{\mathrm{d}}y$ (A.9)
$-\sum_{k\in\Omega\backslash\{-1\}}\frac{f_{k}}{k+1}r^{k+1}-f_{-1}\ln r.$
The second term in (A.9) goes to zero as $r\searrow 0$. After pulling the last two terms to the left-hand side, we may thus perform the limit $r\searrow 0$. We obtain (A.7). $\Box$
Corollary A.5.
Let $f$, $g$ be measurable functions and suppose that $fg^{\prime},gf^{\prime}\in{\cal F}$, and that $\lim_{r\to\infty}f(r)g(r)=0$. Then
$\mathrm{gen}\int_{0}^{\infty}f(r)g^{\prime}(r){\mathrm{d}}r=-\mathrm{gen}\int_{0}^{\infty}f^{\prime}(r)g(r){\mathrm{d}}r-\mathop{\mathrm{rv}_{0}}(fg).$ (A.10)
@@
$f(r)g(r)=-\int_{r}^{\infty}(fg)^{\prime}(y){\mathrm{d}}y.$ (A.11)
Therefore, it is enough to apply (A.7) to $F\coloneqq fg$. $\Box$
A.3 Dimensional regularization
@@
Let $N\in{\mathbb{N}}$ and let $f:\,]0,\infty[\times\{\alpha\in{\mathbb{C}}\,|\,\mathrm{Re}(\alpha)>-N-1\}\to{\mathbb{C}}$ be a function such that $f(r,\cdot)$ is holomorphic for each $r$, $\|f(\cdot,\alpha)\|_{L^{1}[1,\infty[}$ is bounded locally uniformly in $\alpha$, and there exist holomorphic functions $f_{0},\dots,f_{N}$ of $\alpha$ such that the $L^{1}]0,1]$ norm of $f(r,\alpha)-\sum_{n=0}^{N}r^{\alpha+n}f_{n}(\alpha)$ is bounded locally uniformly in $\alpha$. Then $f(\cdot,\alpha)$ is integrable in the generalized sense, and for $-\alpha\not\in\{1,\dots,N\}$ one has
$\quad\mathrm{gen}\int_{0}^{\infty}f(r,\alpha){\mathrm{d}}r$ (A.12)
@@
By Morera’s theorem, the right hand side is, away from the poles at $-1,\dots,-N$, a holomorphic function of $\alpha$. Therefore, to obtain (A.12) in the non-anomalous case it is enough to compute (A.12) in the region where the usual integral is convergent and continue analytically.
Let $m\in\{1,\dots,N\}$. The right hand side of (A.12) has a simple pole at $\alpha=-m$ with residue $f_{m-1}(-m)$ (possibly zero). Its finite part is
$\operatornamewithlimits{fp}_{\alpha\to-m}\mathrm{gen}\int_{0}^{\infty}f(r,\alpha){\mathrm{d}}r=\lim_{\alpha\to-m}\Bigg{(}\mathrm{gen}\int_{0}^{\infty}f(r,\alpha){\mathrm{d}}r-\frac{f_{m-1}(-m)}{\alpha+m}\Bigg{)}.$ (A.13)
@@
$\mathrm{gen}\int_{0}^{\infty}f(r,-m){\mathrm{d}}r=\operatornamewithlimits{fp}_{\alpha\to-m}\mathrm{gen}\int f(r,\alpha){\mathrm{d}}r-f^{\prime}_{m-1}(-m).$ (A.14)
@@ Appendix B Two useful lemmas
$\sum_{k=0}^{n}\frac{(a)_{k}}{k!}=\frac{(a+1)_{n}}{n!},$ (B.1)
which implies for $m,n\in{\mathbb{N}}$, $m>n$:
$\sum_{k=1}^{n+1}\frac{(m-k)!}{(n+1-k)!}=\frac{m!}{n!}\frac{1}{m-n}.$ (B.2)
@@
To prove the first identity, we compute
$\sum_{k=0}^{n}\frac{(a)_{k}}{k!}$ $=\sum_{k=0}^{n}\frac{(a+1)_{k-1}}{k!}(a+k-k)$ (B.3)
$=\sum_{k=0}^{n}\frac{(a+1)_{k}}{k!}-\sum_{k=1}^{n}\frac{(a+1)_{k-1}}{(k-1)!}.$ (B.4)
The second identity follows from the first by setting $a=m-n$. $\Box$
Lemma B.2.
@@
$\frac{\psi^{\prime}(z)}{\Gamma(z)^{2}}\Big{|}_{z=-n}=(n!)^{2},\quad\text{and}\quad\frac{\psi^{\prime}(z)-\psi(z)^{2}}{\Gamma(z)}\Big{|}_{z=-n}$ $=(-1)^{n}2n!\psi(1+n).$ (B.5)
@@
Let $z\in{\mathbb{C}}\setminus-{\mathbb{N}}_{0}$. Then
$\frac{\psi^{\prime}(z)}{\Gamma(z)^{2}}=z^{2}\frac{\psi^{\prime}(z+1)+\frac{1}{z^{2}}}{\Gamma(z+1)^{2}}=z^{2}\frac{\psi^{\prime}(z+1)}{\Gamma(z+1)^{2}}+\frac{1}{\Gamma(z+1)^{2}}.$ (B.6)
Therefore,
$\frac{\psi^{\prime}(z)}{\Gamma(z)^{2}}=\big{(}(z)_{n}\big{)}^{2}\frac{\psi^{\prime}(z+n)}{\Gamma(z+n)^{2}}+\sum_{j=1}^{n}\frac{\big{(}(z)_{j-1}\big{)}^{2}}{\Gamma(z+j)^{2}}.$ (B.7)
Taking the limit $z\to-n$, we obtain
$\frac{\psi^{\prime}(z)}{\Gamma(z)^{2}}\Big{|}_{z=-n}=(n!)^{2}\frac{\psi^{\prime}(z)}{\Gamma(z)^{2}}\Big{|}_{z=0}=(n!)^{2}.$ (B.8)
The proof of the second identity works similar: Using
$\lim_{\epsilon\to 0}\frac{\psi^{\prime}(\epsilon)-\psi(\epsilon)^{2}}{\Gamma(\epsilon)}=-2\gamma_{\rm E}=2\psi(1),$ (B.9)
and
$\frac{\psi^{\prime}(z)-\psi(z)^{2}}{\Gamma(z)}$ $=\frac{\psi^{\prime}(z+1)+\frac{1}{z^{2}}-\big{(}\psi(z+1)-\frac{1}{z}\big{)}^{2}}{\frac{1}{z}\Gamma(1+z)}$ (B.10)
$=z\frac{\psi^{\prime}(z+1)-\psi(z+1)^{2}}{\Gamma(z+1)}+2\frac{\psi(z+1)}{\Gamma(z+1)},$
we inductively obtain
$\frac{\psi^{\prime}(z)-\psi(z)^{2}}{\Gamma(z)}=(z)_{n}\frac{\psi^{\prime}(z+n)-\psi(z+n)^{2}}{\Gamma(z+n)}+2\sum_{j=1}^{n}(z)_{j-1}\frac{\psi(z+j)}{\Gamma(z+j)}.$ (B.11)
Taking the limit $z\to-n$ and using (B.9), we obtain
$\frac{\psi^{\prime}(z)-\psi(z)^{2}}{\Gamma(z)}\Big{|}_{z=-n}$ $=2(-n)_{n}\psi(1)+2\sum_{j=1}^{n}(-n)_{j-1}(-1)^{n-j-1}(n-j)!$ (B.12)
$=(-1)^{n}2n!\Big{(}\psi(1)+\sum_{j=1}^{n}\frac{1}{n+1-j}\Big{)}$
$=(-1)^{n}2n!\psi(1+n).$
$\Box$
Acknowledgement.
The support of the National Science Center of Poland under the grant UMO-2019/35/B/ST1/01651 is acknowledged. J.M.V. was supported by the Academy of Finland grant number 340461 "Conformal invariance in planar random geometry" and the Academy of Finland Centre of Excellence Programme grant number 346315 "Finnish centre of excellence in Randomness and STructures (FiRST)".

### v2

@@ Appendix A Generalized integral
We say that a function $f$ on $0,\infty[$ is integrable in the generalized sense if it is integrable on $1,\infty[$ and there exists a finite set $\Omega\subset{\mathbb{C}}$ and complex coefficients $(f_{k})_{k\in\Omega}$ such that
$f-\sum_{k\in\Omega}f_{k}r^{k}$ (109)
[…]
Note that the set $\{k\in\Omega\,|\,\Re(k)\leq-1\}$ and the corresponding $f_{k}$ are uniquely determined by $f$. It is convenient to allow $k$ with $\Re(k)>-1$. The generalized integral of $f$ does not depend on the choice of $\Omega$.
@@
[…]
@@
The generalized integral (110) is called anomalous if there exists $n\in{\mathbb{N}}$ such that $f_{-n}\neq 0$.
A2. Integration by parts
Definition We say that $F\in{\cal G}$ if $F$ is a measurable function on $0,\infty[$, bounded on $1,\infty[$ and there exists a finite set $\Theta\subset{\mathbb{C}}$ and complex coefficients $(F_{k})_{k\in\Theta}$ and $c\in{\mathbb{C}}$, such that
$F(r)-\sum_{k\in\Theta}F_{k}r^{k}-c\ln r$ (112)
has a limit as $r\to 0$. For $F\in{\cal G}$ we define the regular value of $F$ at 0:
$\mathop{\mathrm{rv}_{0}}F\coloneqq\lim_{r\searrow 0}\Big(F(r)-\sum_{k\in\Theta\backslash\{0\}}F_{k}r^{k}-c\ln r\Big).$ (113)
Note that given $F$ the coefficients $F_{k}$ for $k\in\Theta$ such that $\Re(k)\leq 0,$ $k\neq 0$, as well as $c$, are uniquely defined. Hence (113) depends only on $F$.
@@
$\mathrm{gen}\int_{0}^{\infty}f(y){\mathrm{d}}y=-\mathop{\mathrm{rv}_{0}}F.$ (115)
@@
It is clear that $F$ is bounded on $1,\infty[$ because $f\in{\cal F}$. Let $0<r<1$. Then
$\int_{r}^{\infty}f(y){\rm d}y$ $=\int_{r}^{1}\Big(f(y)-\sum_{k\in\Omega}f_{k}y^{k}\Big){\mathrm{d}}y+\int_{1}^{\infty}f(y){\mathrm{d}}y$ (116)
[…]
$=$ $\mathrm{gen}\int_{0}^{\infty}f(y){\mathrm{d}}y-\int_{0}^{r}\Big(f(y)-\sum_{k\in\Omega}f_{k}y^{k}\Big){\mathrm{d}}y$ (117)
[…]
The second term in (117) goes to zero as $r\searrow 0$. After pulling the last two terms to the left-hand side, we may thus perform the limit $r\searrow 0$. We obtain (115). ∎
et $f$, $g$ be measurable functions and suppose that $fg^{\prime},gf^{\prime}\in{\cal F}$, and that $\lim_{r\to\infty}f(r)g(r)=0$. Then
$\mathrm{gen}\int_{0}^{\infty}f(r)g^{\prime}(r){\mathrm{d}}r=-\mathrm{gen}\int_{0}^{\infty}f^{\prime}(r)g(r){\mathrm{d}}r-\mathop{\mathrm{rv}_{0}}(fg).$ (118)
@@
$f(r)g(r)=-\int_{r}^{\infty}(fg)^{\prime}(y){\mathrm{d}}y.$ (119)
Therefore, it is enough to apply (115) to $F\coloneqq fg$. ∎
A3. Dimensional regularization
@@
Let $N\in{\mathbb{N}}$ and let $f:\,]0,\infty[\times\{\alpha\in{\mathbb{C}}\,|\,\Re(\alpha)>-N-1\}\to{\mathbb{C}}$ be a function such that $f(r,\cdot)$ is holomorphic for each $r$, $\|f(\cdot,\alpha)\|_{L^{1}[1,\infty[}$ is bounded locally uniformly in $\alpha$, and there exist holomorphic functions $f_{0},\dots,f_{N}$ of $\alpha$ such that the $L^{1}]0,1]$ norm of $f(r,\alpha)-\sum_{n=0}^{N}r^{\alpha+n}f_{n}(\alpha)$ is bounded locally uniformly in $\alpha$. Then $f(\cdot,\alpha)$ is integrable in the generalized sense, and for $\alpha\not\in\{1,\dots,N\}$ one has
$\quad\mathrm{gen}\int_{0}^{\infty}f(r,\alpha){\mathrm{d}}r$ (120)
@@
By Morera’s theorem, the right hand side is, away from the poles at $1,\dots,-N$, a holomorphic function of $\alpha$. Therefore, to obtain (120) in the non-anomalous case it is enough to compute (120) in the region where the usual integral is convergent and continue analytically.
Let $m\in\{1,\dots,N\}$. The right hand side of (120) has a simple pole at $\alpha=-m$ with residue $f_{m-1}(-m)$ (possibly zero). Its finite part is
$\operatornamewithlimits{fp}_{\alpha\to-m}\mathrm{gen}\int_{0}^{\infty}f(r,\alpha){\mathrm{d}}r=\lim_{\alpha\to-m}\Bigg(\mathrm{gen}\int_{0}^{\infty}f(r,\alpha){\mathrm{d}}r-\frac{f_{m-1}(-m)}{\alpha+m}\Bigg).$ (121)
@@
$\mathrm{gen}\int_{0}^{\infty}f(r,-m){\mathrm{d}}r=\operatornamewithlimits{fp}_{\alpha\to-m}\mathrm{gen}\int f(r,\alpha){\mathrm{d}}r-f^{\prime}_{m-1}(-m).$ (122)
@@ Appendix B Two useful lemmas
$\sum_{k=0}^{n}\frac{(a)_{k}}{k!}=\frac{(a+1)_{n}}{n!},$ (123)
which implies for $\in{\mathbb{N}}, m>nm>n$:
$\sum_{k=1}^{n+1}\frac{(m-k)!}{(n+1-k)!}=\frac{m!}{n!}\frac{1}{m-n}.$ (124)
@@
The first identity follows from a standard telescoping sum argument. The second identity follows from the first by setting a=mna=m-n. ∎
@@
$\frac{\psi^{\prime}(z)}{\Gamma(z)^{2}}\Big|_{z=-n}=(n!)^{2},\quad\text{and}\quad\frac{\psi^{\prime}(z)-\psi(z)^{2}}{\Gamma(z)}\Big|_{z=-n}$ $=(-1)^{n}2n!\psi(1+n).$ (125)
@@
Let $z\in{\mathbb{C}}\setminus-{\mathbb{N}}_{0}$. It is easy to verify that for $n\in{\mathbb{N}}$, we have
$\frac{\psi^{\prime}(z)}{\Gamma(z)^{2}}=\big((z)_{n}\big)^{2}\frac{\psi^{\prime}(z+n)}{\Gamma(z+n)^{2}}+\sum_{j=1}^{n}\frac{\big((z)_{j-1}\big)^{2}}{\Gamma(z+j)^{2}}.$ (126)
The first identity now follows from taking the limit $z\to-n$. The proof of the second identity works similarly. We first check that
$\frac{\psi^{\prime}(z)-\psi(z)^{2}}{\Gamma(z)}$ $=(z)_{n}\frac{\psi^{\prime}(z+n)-\psi(z+n)^{2}}{\Gamma(z+n)}+2\sum_{j=1}^{n}(z)_{j-1}\frac{\psi(z+j)}{\Gamma(z+j)}.$ (127)
The claim then follows by taking $z\to-n$, using
$\lim_{\epsilon\to 0}\frac{\psi^{\prime}(\epsilon)-\psi(\epsilon)^{2}}{\Gamma(\epsilon)}=2\psi(1)$ (128)
and standard identities satisfied by $\psi(z)$. ∎